A sailor strikes the side of his ship just below the waterline. He hears the echo of the sound reflected from the ocean floor directly below 2.0 s later.

How deep is the ocean at this point? Assume the speed of sound in seawater is 1560 m/s and does not vary significantly with depth.
Answer #1
let the depth of the ocean be h
total distance travel by sound S = 2h
time of echo t = 2 s
speed of sound in sea water v = 1560 m / s
we know v = S / t = 2h / t
from this h = vt / 2
= 1560 m